# A Plea for History and Philosophy of Statistics and Machine Learning

### v1

@@ Abstract
Abstract: The integration of the history and philosophy of statistics was initiated at least by Hacking (1965) and advanced by Mayo (1996), but it has not received sustained follow-up. Yet such integration is more urgent than ever, as the recent success of artificial intelligence has been driven largely by machine learning—a field historically developed alongside statistics. Today, the boundary between statistics and machine learning is increasingly blurred. What we now need is integration, twice over: of history and philosophy, and of the field they engage—statistics and machine learning. I present a case study of a philosophical idea in machine learning (and in formal epistemology) whose root can be traced back to an often under-appreciated insight in Neyman and Pearson’s 1936 work (a follow-up to their 1933 classic). This leads to the articulation of an epistemological principle—largely implicit in, but shared by, the practices of frequentist statistics and machine learning—which I call achievabilism: the thesis that the correct standard for assessing non-deductive inference methods should not be fixed, but should instead be sensitive to what is achievable in specific problem contexts. Another integration also emerges at the level of methodology, combining two ends of the philosophy of science spectrum: history and philosophy of science on the one hand, and formal epistemology on the other hand.
@@ 1 Introduction
  • Descriptive Question: Descriptive Question: Are there interesting epistemological views about scientific inference that remain largely implicit in the practices of statistics, machine learning, or other fields that study scientific inference (such as econometrics), and that still need to be articulated and made explicit?
@@ 2.1 Physical Probabilities
While this interpretational issue is important, it must be set aside here. My primary task is to articulate certain epistemological principles implicit in the practice of frequentist statistics. I want to acknowledge their reference to physical probabilities, but must leave the interpretation of those probabilities open due to space constraints. Suffice it to note that the probabilities used in frequentist statistics might—or might not—be best understood as long-run relative frequencies, since notable alternative interpretations exist, as mentioned above. The term ‘frequentist’ in ‘frequentist statistics’ is therefore quite misleading—but unfortunately too entrenched to change.
@@ 2.4 Inductive Risks
2.5 Risk-World Plots
@@
Inductive risk can be plotted not only against possible worlds $\theta$ with a fixed sample size $n$, but also against pairs $(\theta,n)$, where both the possible world $\theta$ and the sample size $n$ vary. The resulting graph is a “surface” rising above a two-“dimensional plane” spanned by the axis of possible worlds and the axis of sample sizes.
@@ 4 Negligence in Statistics and in Philosophy
Worse, achievabilism continued to be overlooked by Lehmann, a student of Neyman’s and author of the influential, now-classic textbook Testing Statistical Hypotheses. The first edition (1959) did motivate the lower standard of UMPU on page 125, presenting it as the result of lowering the bar from UMP, closely following Neyman and Pearson’s (1936) rationale. A true achievabilist would then immediately ask: now that the bar has already been lowered once, might there be a problem context in which UMPU itself is unachievable, requiring a further lowering of the bar? Yet this natural question is not raised until much later—over a hundred pages on, at page 228. Note that I am not even talking about addressing the question, but merely about acknowledging it. The fourth edition (2022) shows no improvement on this front, retaining the same long delay. The spirit of achievabilism still seems to be taken lightly.
@@ 5 Repeated Reinventions of Achievabilism
The idea of achievabilism was reinvented at least twice in theoretical computer science: first, in the 1960s, within computability theory and algorithmic learning theory—an approach that eventually made its way into a branch of philosophy, formal epistemology; and then again in the 1980s, in the development of statistical learning theory, which has become one of the most influential approaches to the foundations of machine learning.
@@ 5.1 Curve Fitting: Frequentist Statistics, 1980s-1990s
We are now confronted with the task of assessing AIC and BIC as two competing inference methods. Which version of Ockham’s razor should we adopt—and in what contexts? A somewhat achievabilist answer began to emerge from Shibata’s work in the 1980s. But before introducing it, another brief tutorial is in order.
@@ 5.1.2 A Tutorial on Regression via Point Estimation
Clarification on terminology: While the goal in the present context is to find a good predictive model, the term ‘consistency’ is also often used relative to a different goal: that of identifying the true model. This alternative notion is known as model selection consistency, and it will not concern us in this subsection. The uses of the term ‘consistency’ in statistics (and machine learning) can be quite confusing; context matters. End of this tutorial.
@@
So, an all-too-obvious extension of the above hierarchy is the following:
@@ 5.2 Classification: Machine Learning, 1970s-2010s
Please write a 500-word essay on C. S. Peirce and his view on scientific inference. [End of Request] [Start of Response] Here you go: Charles Sanders Peirce is a philosopher known
@@
While a classifier is itself an algorithm, computer scientists are also interested in a different kind of algorithm: one that outputs a classifier when given a training dataset—a set of annotated images labeled with 1’ or 0’ by human annotators, much like a scatterplot on the $XY$ “plane”). Such an algorithm is called a learning algorithm—an algorithm of a higher order, since its output, a classifier, is itself an algorithm. Learning algorithms for classification in machine learning are thus the counterparts of curve fitting methods for regression in statistics. As a result, most evaluative standards from one area can be readily carried over to the other.
Interestingly, while the high standard of uniform consistency ought to be an obvious object of study, it has drawn attention primarily in the field of classification—and largely among theoretical computer scientists rather than statisticians. In a 1982 paper, the computer scientist Devroye proved that no learning algorithm achieves uniform consistency in problem contexts where the background assumption is weak (imposing almost no constraint on the true, unknown probability distribution over the possible images of cats or non-cats). This impossibility result for classification can be easily carried over to regression.
@@
  • In a context where the set of competing classifiers is relatively small (in the technical sense of having a finite VC dimension), uniform consistency is achievable (Vapnik & Chervonenkis 1974).
  • In a context where the set of competing classifiers is relatively large (with an infinite VC dimension), uniform consistency is unachievable (Devroye 1982) but consistency is achievable (Lugosi & Zeger 1996).
Coincidentally, 1996 saw more than one monograph that painted a distinctly achievabilist picture. One came from machine learning: Devroye and colleagues’ book. Another came from philosophy: Kelly’s The Logic of Reliable Inquiry (1996), whose achievabilist heritage can be traced back to Putnam’s (1965) work in theoretical computer science—but that is an arc of the story in philosophy I will tell later.
@@
Readers of that 2014 textbook can easily reconstruct a hierarchy of standards for assessing learning algorithms:
@@
The exact definition of the intermediate standard—nonuniform learnability—is not important for now. What matters is this: in problem contexts where the set of competing classifiers is large, we already know from Devroye’s (1982) no-free-lunch theorem that the high standard of uniform consistency is unachievable. We also know from Lugosi & Zeger (1996) that the low standard of consistency is achievable; in fact, they proved that it is achieved by a version of Ockham’s razor for model selection in classification, known as structural risk minimization (SRM).
@@ 5.3 Inductive Inference: Theoretical Computer Science and Formal Epistemology, 1960s-1990s
Good news: almost the same idea was developed by the computer scientist Gold (1967), who applied it to evaluate algorithms aimed at explicitly learning something about the empirical world—specifically, learning the grammar of a language by observing utterances from native speakers. Although Gold studied only two standards in Putnam’s hierarchy—the highest and the lowest—his work inspired a generation of researchers who embraced the achievabilist spirit and went on to explore intermediate standards as well. [^5] See Jain & Stephan (2017) and Case & Jain (2017) for recent surveys.
@@ 6.1 Against Norton Against Formal Theories of Inference
In short: formality, when properly construed, does not imply universality. A theory can be formal and yet context-sensitive—and achievabilism is a clear example of this, an example that grew out of the actual practice of scientists who studied scientific inference.
To be sure, Norton’s implication might be saved by restricting its scope—by strengthening the antecedent to invoke a narrower sense of formality. I suspect that, in that case, achievabilism would no longer be categorized as a “formal” theory of scientific inference. Yet achievabilism still plays an important role here: to help us sharpen Norton’s implication and identify a plausible version of it.
@@ 6.2 Unified Justification of Ockham’s Many Local Razors
    Ockham’s (blunt) AIC razor vs. his (sharp) BIC razor in contexts of regression and model selection (Section 5.1);
@@ 6.3 The Possibility of Internalist Reliabilism
According to the conventional version of reliabilism (Goldman 1979), the justification of an inference method $M$ depends solely on the actual reliability that $M$ possesses. This actual reliability—or unreliability—may be measured by the actual inductive risk of $M$: the inductive risk that $M$ has at the actual world. This notion of inductive risk is external in the sense that it is not always possible, even in principle, for an agent to come to recognize the actual inductive risk of their method through mere reflection or introspection. And it is in this sense that conventional reliabilism is undeniably externalist.
In contrast, the implementations of achievabilism across statistics, machine learning, and formal epistemology are internalist. The crucial reason is that the evaluative standards under consideration—such as uniform consistency—do not really rely on the inductive risk of an inference method at the actual world. Instead, they refer to the inductive risk of the assessed method at each possible world compatible with the background assumptions that one makes in one’s context of inquiry. This is something that one can, in principle, examine by mere reflection or introspection alone. Let me highlight the elements that make this approach internalist:
When an evaluative standard refers to inductive risk, it is always the inductive risk${}_{\,(c)}$ of an inference method at a possible world compatible${}_{\,(b)}$ with one’s background ${}_{\,(a)}$, where the three elements $(a)$-$(c)$ are all internalist:
@@
While the implementations of achievabilism across statistics, machine learning, and formal epistemology are clearly internalist, as seen above, they are reliabilist nonetheless—in two important senses. First and foremost, they are reliabilist in that they assign a central role to inductive risk when defining evaluative standards. Second, they are reliabilist in taking the pursuit of reliability seriously: they strive to enforce the highest achievable standard for minimizing, controlling, or diminishing inductive risk across a range of (internally accessible) possible worlds.
@@ 7 Closing
Open Question II (from Machine Learning). In the section on classification, I discussed the development of SLT (statistical learning theory) as an instance of achievabilist thinking—an approach that has long aspired to serve as the (frequentist) epistemological foundation of machine learning. However, it has recently been argued—most notably by computer scientists such as Moritz Hardt—that SLT has little to do with the actual advancement of one of the most fascinating and successful branches of modern machine learning: deep learning. In his keynote speech at the 2024 ICLR (International Conference on Learning Representations), Hardt proposed that deep learning requires a new epistemological foundation, which he dubbed the emerging science of benchmarks (https://iclr.cc/virtual/2024/invited-talk/21799). Is statistical learning theory mistaken? Does the emerging science of benchmarks complement, or compete with, statistical learning theory?

### v2

@@ Abstract
Abstract: The integration of the history and philosophy of statistics was initiated at least by Hacking (1965) and advanced by Mayo (1996), but it has not received sustained follow-up. Yet such integration is more urgent than ever, as the recent success of artificial intelligence has been driven largely by machine learning—a field historically developed alongside statistics. Today, the boundary between statistics and machine learning is increasingly blurred. What we now need is integration, twice over: of history and philosophy, and of two fields they engage—statistics and machine learning. I present a case study of a philosophical idea in machine learning (and in formal epistemology) whose root can be traced back to an often under-appreciated insight in Neyman and Pearson’s 1936 work (a follow-up to their 1933 classic). This leads to the articulation of an epistemological principle—largely implicit in, but shared by, the practices of frequentist statistics and machine learning—which I call achievabilism: the thesis that the correct standard for assessing non-deductive inference methods should not be fixed, but should instead be sensitive to what is achievable in specific problem contexts. Another integration also emerges at the level of methodology, combining two ends of the philosophy of science spectrum: history and philosophy of science on the one hand, and formal epistemology on the other hand.
@@ 1 Introduction
  • Descriptive Question: Are there interesting epistemological views about scientific inference that remain largely implicit in the practices of statistics, machine learning, or other fields that study scientific inference (such as econometrics), and that still need to be articulated and made explicit?
@@ 2.1 Physical Probabilities
At this point, you might be wondering what physical probabilities are, or how the term ‘physical probability’ should be interpreted. This is an important metaphysical—or semantic—question. Perhaps physical probabilities are long-run relative frequencies (Neyman 1955, von Mises 1957), propensities (Popper 1959), Humean chances (Lewis 1981) or primitive physical states posited in science (Sober 2000: sec. 3.2). For focus, I will set aside this metaphysical/semantic debate and concentrate on epistemological issues. Suffice it to note that the probabilities used in frequentist statistics might—or might not—be best understood as long-run relative frequencies, since there are notable alternative interpretations. The term ‘frequentist’ in ‘frequentist statistics’ is therefore quite misleading—but unfortunately too entrenched to change.
@@ 2.4 Inductive Risk
2.5 Risk-World Plot
@@
Inductive risk can be plotted not only against possible worlds $\theta$ with a fixed sample size $n$, but also against pairs $(\theta,n)$, where both the possible world $\theta$ and the sample size $n$ vary. The resulting graph is a “surface” rising above a “two-dimensional plane” spanned by the axis of possible worlds and the axis of sample sizes.
@@ 4 Negligence in Statistics and in Philosophy
Worse, achievabilism continued to be overlooked by Lehmann, a student of Neyman’s and author of the influential, now-classic textbook Testing Statistical Hypotheses. The first edition (1959) did motivate the lower standard of UMPU on page 125, presenting it as the result of lowering the bar from UMP, closely following Neyman and Pearson’s (1936) rationale. A true achievabilist would then immediately ask: now that the bar has already been lowered once, might there be a problem context in which UMPU itself is unachievable, requiring a further lowering of the bar? Yet this natural question is not raised until much later—over a hundred pages on, at page 228. Note that I am not even talking about addressing the question, but merely about acknowledging it. The fourth edition (2022) shows no improvement on this front, retaining the same long delay. Here, the spirit of achievabilism seems to be taken lightly.
@@ 5 Repeated Reinventions of Achievabilism
The idea of achievabilism was reinvented at least twice in theoretical computer science. The first reinvention occurred in the 1960s, within computability theory and algorithmic learning theory, and eventually made its way into a branch of philosophy known as formal epistemology. The second came in the 1980s, with the development of statistical learning theory, which has become one of the most influential approaches to the foundations of machine learning.
@@ 5.1 Curve Fitting: Frequentist Statistics, 1980s-1990s
We are now confronted with the task of assessing AIC and BIC as two competing inference methods. Which version of Ockham’s razor should we adopt—and in what contexts? A somewhat achievabilist answer began to emerge from the statistician Shibata’s work in the 1980s. But before introducing it, another brief tutorial is in order.
@@ 5.1.2 A Tutorial on Regression via Point Estimation
Clarification on terminology: While the goal in the present context is to find a good predictive model, the term ‘consistency’ is also often used relative to a different goal: that of identifying the true model. This alternative notion is known as model selection consistency, and it will not concern us in this section. The uses of the term ‘consistency’ in statistics (and machine learning) can be quite confusing; context matters. End of this tutorial.
@@
So, an obvious extension of the above hierarchy is the following:
@@ 5.2 Classification: Machine Learning, 1970s-2010s
Please write a 500-word essay on C. S. Peirce and his view on
scientific inference. [End of Request] [Start of Response]
Charles Sanders Peirce is a philosopher known
@@
While a classifier is itself an algorithm, computer scientists are also interested in a different kind of algorithm: one that outputs a classifier when given a training dataset—a set of annotated images labeled with ‘1’ or ‘0’ by human annotators, much like a scatterplot on the $XY$ “plane”. Such an algorithm is called a learning algorithm—an algorithm of a higher order, since its output, a classifier, is itself an algorithm. Learning algorithms for classification in machine learning are thus the counterparts of curve fitting methods for regression in statistics. As a result, evaluative standards from one area can often be carried over to the other.
Interestingly, while the high standard of uniform consistency ought to be an obvious object of study, it has drawn attention primarily in the field of classification—and largely among theoretical computer scientists rather than statisticians. In a 1982 paper, the computer scientist Devroye proved that no learning algorithm achieves uniform consistency in problem contexts of classification where the background assumption is weak (imposing almost no constraint on the true, unknown probability distribution over the possible images of cats or non-cats). This impossibility result for classification can be easily carried over to regression.
@@
  • In a context where the set of competing classifiers to choose from is relatively small (in the technical sense of having a finite VC dimension), uniform consistency is achievable (Vapnik & Chervonenkis 1974).
  • In a context where the set of competing classifiers to choose from is relatively large (with an infinite VC dimension), uniform consistency is unachievable (Devroye 1982) but consistency is achievable (Lugosi & Zeger 1996).
Coincidentally, 1996 saw more than one monograph that painted a distinctly achievabilist picture. One came from machine learning: Devroye and colleagues’ book, as we have seen. Another came from philosophy: Kelly’s The Logic of Reliable Inquiry (1996), whose achievabilist heritage can be traced back to Putnam’s (1965) work in theoretical computer science—but that is an arc of the story in philosophy I will tell later.
@@
Readers of Shalev-Shwartz & Ben-David’s 2014 textbook can easily reconstruct a hierarchy of standards for assessing learning algorithms:
@@
(Terminological note: Shalev-Shwartz & Ben-David uses the more descriptive term ‘convergence’ instead of the less intuitive term ‘consistency’.) The exact definition of the intermediate standard—nonuniform learnability—is not important for now. What matters is this: in problem contexts where the set of competing classifiers is large, we already know from Devroye’s (1982) no-free-lunch theorem that the high standard of uniform consistency is unachievable. Yet the low standard of consistency is achievable in those problem contexts, thanks to Lugosi & Zeger (1996); in fact, they proved that it is achieved by a version of Ockham’s razor for model selection in classification, known as structural risk minimization (SRM).
@@ 5.3 Inductive Inference: Theoretical Computer Science and Formal Epistemology, 1960s-1990s
Good news: almost the same idea was developed by the computer scientist Gold (1967), who applied it to evaluate algorithms aimed at explicitly learning something about the empirical world—specifically, learning the grammar of a language by observing utterances from native speakers. Although Gold studied only two standards in Putnam’s hierarchy—the highest and the lowest—his work inspired a generation of theoretical computer scientists who embraced the achievabilist spirit and went on to explore intermediate standards as well. For recent surveys, see Jain & Stephan (2017) and Case & Jain (2017).
@@ 6.1 Against Norton Against Formal Theories of Inference
In short: formality, when properly construed, does not imply universality. A theory can be formal and yet context-sensitive, and achievabilism is a clear example of this—an example that grew out of the actual practice of scientists who studied scientific inference.
To be sure, Norton’s implication might be saved by restricting its scope—by strengthening the antecedent to invoke a narrower sense of formality. I suspect that, in that case, achievabilism would no longer be categorized as a “formal” theory of scientific inference, but it should still play an important role: to help us sharpen Norton’s implication and identify a plausible version of it.
@@ 6.2 Unified Justification of Ockham’s Many Local Razors
    Ockham’s (blunt) AIC razor vs. (sharp) BIC razor in contexts of regression and model selection (Section 5.1);
@@ 6.3 The Possibility of Internalist Reliabilism
According to the conventional version of reliabilism (Goldman 1979), the justification of an inference method $M$ depends solely on the actual reliability that $M$ possesses. This actual reliability—or unreliability—may be measured by the actual inductive risk of $M$: the inductive risk that $M$ has at the actual world. This factor of justification—actual inductive risk—is external in the sense that it is not always possible, even in principle, for an agent to come to recognize the actual inductive risk of their method through mere reflection or introspection. And it is in this sense that conventional reliabilism is undeniably externalist.
In contrast, the implementations of achievabilism across statistics, machine learning, and formal epistemology are internalist. The crucial reason is that the evaluative standards under consideration—such as uniform consistency—do not really rely on the inductive risks at the actual world. Instead, they refer to the inductive risk of the assessed method at each of the possible worlds compatible with the background assumptions that one makes in one’s context of inquiry—whether or not the actual world turns out to be one of those worlds. Such inductive risks are something that one can, in principle, examine by mere reflection or introspection alone. Let me highlight the elements that make this approach internalist:
When an evaluative standard refers to inductive risk, it is always the inductive risk${}_{\,(c)}$ of an inference method at a possible world compatible${}_{\,(b)}$ with one’s background ${}_{\,(a)}$, where the three underlined elements $(a)$-$(c)$ are all internalist:
@@
While the implementations of achievabilism across statistics, machine learning, and formal epistemology are clearly internalist, as seen above, they are reliabilist nonetheless—in two important senses. First, they are reliabilist in that they assign a central role to inductive risk when defining evaluative standards. Second, they are reliabilist in taking the pursuit of reliability seriously: they strive to enforce the highest achievable standard for minimizing, controlling, or diminishing inductive risk across a range of (internally accessible) possible worlds.
@@ 7 Closing
Open Question II (from Machine Learning). In the section on classification, I discussed the development of statistical learning theory as an instance of achievabilist thinking—an approach that has long aspired to serve as the (frequentist) epistemological foundation of machine learning. However, it has recently been argued—most notably by computer scientists such as Moritz Hardt—that statistical learning theory has little to do with the actual advancement of one of the most fascinating and successful branches of modern machine learning: deep learning. In his keynote speech at the 2024 ICLR (International Conference on Learning Representations), Hardt proposed that deep learning requires a new epistemological foundation, which he dubbed the emerging science of benchmarks (https://iclr.cc/virtual/2024/invited-talk/21799). Is statistical learning theory mistaken? Does the emerging science of benchmarks complement, or compete with, statistical learning theory?